327 mg acety c acid (HC H-04). Calculate the pH of a A typical aspirin tablet in enough water to make one cup (237 mL.) of solution. Assume the aspirin tablets are pure c acid, K3.3x 10-4 pH
Answer #1

mass of aspirin = 327 mg = 0.327 g

molar mass of aspirin = 180 g / mol

moles of aspirin = 0.327 / 180

                           = 1.82 x 10^-3

volume = 237 mL = 0.237 L

molarity = 1.82 x 10^-3 / 0.237 = 7.67 x 10^-3

             = 0.00767 M

HA ----------------------> H+ + A-

0.00767                       0         0

-x                              + x          +x

0.00767-x                  x           x

Ka = x^2 / 0.00767 -x = 3.3 x 10^-4

x^2 + 3.3 x 10^-4 x - 2.53 x 10^-6

x = [H+] = 1.43 x 10^-3 M

pH = -log[ H+ ]

pH = 2.84